When you retire 50 years from now, you want to have $2,000,000 million. You think you can earn an average of 8 percent on your investments. To meet your goal, you are trying to decide whether to deposit a lump sum today, or to wait and deposit a lump sum 2 years from today. How much more will you have to deposit as a lump sum if you wait for 2 years before making the deposit? Group of answer choices $3,280 $4,107 $5,289 $6,188 $7,096

What is the present value of the following set of end of the year cash flows: Year 1: $450; Year 2: $600; and then an eight year annuity of $400 per year if the discount rate is 9%?

$2,214 $2,462 $2,781 $3,112 $3,244
Answer #1

Amount Requird Today :

FV = PV (1+r)^n

2,000,000 = PV (1+0.08)^50

2,000,000 = PV (1.08)^50

2,000,000 = PV * 46.9016

PV = 2,000,000 / 46.9016

= 42642.46

Amount Requird after 2 Years :

2,000,000 = PV (1+0.08)^48

2,000,000 = PV (1.08)^48

2,000,000 = PV * 40.2106

PV = 2,000,000 / 40.2106

= 49738.16

Extra needed = 49738.16 - 42642.46

= 7095.70 i.e 7096.

Present Value:

DIsc CF = CF * PVF(r%, n)

PVF = 1 / (1+r)^n

Year CF PVF @9% Disc CF
1 $ 450.00     0.9174 $    412.84
2 $ 600.00     0.8417 $    505.01
3 $ 400.00     0.7722 $    308.87
4 $ 400.00     0.7084 $    283.37
5 $ 400.00     0.6499 $    259.97
6 $ 400.00     0.5963 $    238.51
7 $ 400.00     0.5470 $    218.81
8 $ 400.00     0.5019 $    200.75
9 $ 400.00     0.4604 $    184.17
10 $ 400.00     0.4224 $    168.96
PV of Cash flows $ 2,781.27